4.At a zoo, a sampling of children was asked if the zoo were to get one additional animal, would they prefer a lion or an elephant. The results of the survey follow:

Write each of the following in terms of events (L/E/B/G, respectively; for example if asked what the probability that a selected child prefers a Lion, given that the child is a Boy, you would

respond: P(L|B) = 90 ) 200

What is the probability that:
a. A randomly selected child is a Boy?
b. A randomly selected Girl prefers Elephants?
c. A randomly selected child that prefers Lions is a Boy d. A randomly selected child prefers Elephants?

(please I need complete answers)
Answer #1

a)P( a randomly selected child is a boy) = P(B)

b) P( a randomly selected girl prefer elephants)=P(E|B)

c) P(A randomly selected child that prefers Lions is a Boy) = P( B| L)

d) P( a randomly selected child prefer elephants) = P(E)

I hope u are looking for this notations only as any other information is not given, so i cannot count the probability.